Consider an estimated population of 602 million people.

(a) What is the maximum number of IP addresses that can be assigned per person using IPv4 addressing?

(b) Design an appropriate CIDR to deliver a host IP address for each person of this population
Answer #1
1)
number of bits in an IP address is 32
so, total number of IP addresses is 2^32 = 4294967296
population = 602,000,000
number of IP addresses per person = 4294967296/602000000 = 7
Answer: 7

2)
CIDR is /28
because 2^28 is 268435456